OpenStudy (anonymous):

An initial investment of $1000 is appreciated for 8 years in an account that earns 9% interest, compounded annually. Find the amount of money in the account at the end of the period. Let f(x) = 3x. Find f( 2). Solve the equation 1296 x = 6

OpenStudy (anonymous):

These are three different questions. I will solve eat one in a separated post.

OpenStudy (anonymous):

The amount of money by the end of the 8th year is \(1000(1+0.09)^8=$1993\).

OpenStudy (anonymous):

For the second question, just plug \(x=2\) to find f(2). That's \(f(2)=3(2)=6\).

OpenStudy (anonymous):

In the last problem, divide both sides by \(1296\) to get x. \(x=\frac{6}{1296}=\frac{1}{216}\).

OpenStudy (anonymous):

3 separate questions? 1: annually compounding interest formula FV=P(1+i/r)^(rt) FV = Future Value P = Present Value i = interest rate r = compounding frequency (annually, monthly, so on) t = time (interest periods how many periods pass so, in your case: FV=1000(1+(.09))^(8) FV=1992.56 2. f(2)=6 because you substitute 2 in for x and 2*3 is six 3. x=6/1296

OpenStudy (anonymous):

yeah for questions 2 and 3, I thought to answer them so but wasnt sure if they had anything to do with question 1

OpenStudy (anonymous):

thanks!